. For a pre-lunch snack, Cristal consumes 18 crackers from a box of Tajablanks crackers with the following on the information panel:

            Serving size 9 crackers (0.42 oz)       Fat 0.222 oz (per serving size)

            Protein 0.099 oz (per serving size)     Carbohydrate 0.099 oz (per serving size)

What is the energy in kilojoules that is due to the fat in Cristal’s snack?
Answer #1

Answer-

Given,

Crackers consumed = 18

Serving Size = 9 crackers

Mass of each Servings = 0.42 oz

Mass of Fat in Each serving = 0.222 oz

Mass of Protein in each Serving = 0.099 oz

Mass of Carbohydrate in each servings = 0.099 oz

Energy in Kilo Joule due to fat = ?

We know that,

Serving size = 9 crackers

Total Crackers consumed = 18

So, No. of servings consumed = 18/9 = 2

fat in one serving = 0.222 oz

So, Total Fat consumed = 2 * 0.222 oz = 0.444 oz

1 oz = 28.3495 g

So, 0.444 oz = 12.59 g

Also,

Energy due Fat = 37 kJ/g

So, Energy due to 0.444 oz or 12.59 g fat = 37 kJ/g * 12.59 g

Energy due to 0.444 oz or 12.59 g fat = 465.83 kJ [Answer]